You hold a diversified $100,000 portfolio consisting of 20 stocks with $5,000 invested in each. The portfolio's beta is 1.12. You plan to sell a stock with b = 0.90 and use the proceeds to buy a new stock with b = 1.50. What will the portfolio's new beta be? Do not round your intermediate calculations.
    a. 1.185  
   b. 1.242  
   c. 1.127  
   d. 1.150  
   e. 1.093
Answer #1

Increase in beta as a result of net sales=(1.5-0.9)

=0.6

Hence increase/stock=0.6/20

=0.03

Hence new beta=1.12+0.03

=1.15